# The $S_{n}$-equivariant Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$

### v1

@@ 3.4. Separating rational tails
The term $[\operatorname{Spec}\mathbb{C}]$ in the formula has bidegree $n=1,d=0$ and is interpreted as providing legs on the core. In the following proof, it will be convenient to interpret the term by formally setting ¯$\overline{\mathcal{M}}_{0}^{\star}(r)(1,0):=\operatorname{Spec}\mathbb{C}$, whereas ¯$\overline{\mathcal{M}}_{0}^{\star}(r)(0,0)=\varnothing$ because of the stability condition.
@@
[…]
@@
Denote the item in the above formula indexed by $\delta,j$ as $X_{\delta,j}$. We claim that $X_{\delta,j}$ has the same class as the locally closed stratum ¯$\overline{\mathcal{M}}^{(\delta,j)}_{1,n}(\mathbb{P}^{r},d)$ in ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ consisting of stable maps where the core has degree $\delta$, and the core has valency $j$: in other words, the number of rational tails and the number of legs on the core sum to $j$.
To prove the claim, we further stratify ¯$\overline{\mathcal{M}}^{(\delta,j)}_{1,n}(\mathbb{P}^{r},d)$ by the unordered partitions $\lambda\vdash n$ and $\mu\vdash(d-\delta)$ into $j$ non-negative parts, which record the distribution of the number of marked points and degrees onto each leg or rational tail. Let ¯$\overline{\mathcal{M}}^{(\lambda,\mu)}_{1,n}(\mathbb{P}^{r},d)$ be such a refined stratum. From Lemmas 3.6, 3.7, we have that
¯$\overline{\mathcal{M}}^{(\lambda,\mu)}_{1,n}(\mathbb{P}^{r},d)=\operatorname{Ind}^{S_{n}}_{S_{\lambda}}\left[\overline{\mathcal{M}}^{\mathrm{nrt}}_{1,j}(\mathbb{P}^{r},\delta)\times\prod_{i=1}^{j}\overline{\mathcal{M}}^{\star}_{0}(r)(\lambda_{i},\mu_{i})\right]/\mathrm{Aut}(\lambda,\mu),$
@@
where each $Y_{(\lambda,\mu)}$ consists of union of $j$-fold products in $\left(\mathsf{pt}_{S_{1}}+\overline{\mathcal{M}}_{0}^{*}(r)\right)^{\boxtimes j}(n,d-\delta)$ such that the unordered partitions on marked points and degrees recover $\lambda,\mu$. More explicitly, consider the set $L_{(\lambda,\mu)}$ of pairs of ordered partitions that forget to $(\lambda,\mu)$, then $Y_{(\lambda,\mu)}$ is a disjoint
@@
Applying Lemma 3.1 to the $S_{j}$-action on ¯$\overline{\mathcal{M}}^{\mathrm{nrt}}_{1,j}(\mathbb{P}^{r},\delta)\times Y_{(\lambda,\mu)}$, we hence identify
[¯$\left[\overline{\mathcal{M}}_{1,j}^{\mathrm{nrt}}(\mathbb{P}^{r},\delta)\times Y_{(\lambda,\mu)}\right]/S_{j}\cong\overline{\mathcal{M}}^{(\lambda,\mu)}_{1,n}(\mathbb{P}^{r},d).$
Notice that all the equalities and isomorphisms above are $S_{n}$-equivariant. Summing over all partitions $(\lambda,\nu)$ and all $(\delta,j)$, we get the desired formula in the equivariant Grothendieck ring. ∎
@@ 4. Localisation on the Kontsevich space
[…]
@@
In this section, we introduce a combinatorial graded $\mathbb{B}$-space $\mathsf{Dih}_{r}$ and an $(S_{2}\times\mathbb{S})$-space $\mathsf{Cat}$ of chains of rational curves, and prove the following theorem for the motive [¯$[\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)]\in K_{0}(\mathsf{V},\mathbb{S})[[q]]$.
@@
[¯$[\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)^{\mathbb{C}^{\star}}]=[({\mathbb{P}^{r}})^{\mathbb{C}^{\star}}][\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(0)]+[\mathsf{Dih}_{r}]\circ_{S_{2}}[\mathsf{Cat}].$
Graber and Pandharipande [GP99, §4] have described the components
¯$\overline{\mathcal{M}}_{g}(r)^{\mathbb{C}^{\star}}(n,d)=\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$
of the graded $\mathbb{S}$-space ¯$\overline{\mathcal{M}}_{g}(r)^{\mathbb{C}^{\star}}$ in terms of certain decorated graphs, as we recall shortly. Theorem 4.1 is a motivic interpretation of their description when $g=1$.
@@ 4.1. Localisation graphs
is a graph colouring of $\mathbf{G}$ with $r+1$ colours. The isomorphisms in $\mathsf{LG}_{g,n,d}(r)$ are $\mathsf{LG}_{g,n,d}$-isomorphisms that also respect the colouring. For convenience, we set $\mathsf{LG}_{g,n,d}(0):=\mathsf{LG}_{g,n,d}$. An object $(\mathbf{G},c)$ of the groupoid $\mathsf{LG}_{g,n,d}(r)$ corresponds to a $\mathbb{C}^{\star}$-fixed closed subscheme ¯$\overline{\mathcal{M}}_{\mathbf{G},c}\subset\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ with
(4.2) ¯$\overline{\mathcal{M}}_{\mathbf{G},c}\cong\prod_{v\in V(\mathbf{G})}\overline{\mathcal{M}}_{w(v),n(v)+\mathrm{val}(v)}/\mathrm{Aut}_{\mathsf{LG}_{g,n,d}(r)}(\mathbf{G},c),$
@@
(4.3) ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}=\coprod_{(\mathbf{G},c)\in\pi_{0}(\mathsf{LG}_{g,n,d}(r))}\overline{\mathcal{M}}_{\mathbf{G},c},$
@@
Then $\operatorname{Aut}(\mathcal{G},c)$ is a subgroup of $\operatorname{Aut}(\mathcal{G})$. Each $(\mathcal{G},c)$ in the groupoid $\widetilde{\mathsf{LG}}_{g,n,d}(r)$ defines an $S_{n}$-equivariant $\mathbb{C}^{\star}$-fixed subscheme ¯$\overline{\mathcal{M}}_{\mathcal{G},c}$ of ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$, leading to $S_{n}$-equivariant versions of equations (4.2) and (4.3):
(4.4) ¯$\overline{\mathcal{M}}_{\mathcal{G},c}\cong\left(\operatorname{Ind}_{\prod_{v\in V(\mathcal{G})}S_{n(v)}}^{S_{n}}\prod_{v\in V(\mathcal{G})}\overline{\mathcal{M}}_{g(v),n(v)+\operatorname{val}(v)}\right)/\operatorname{Aut}(\mathcal{G},c),$
(4.5) ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}=\coprod_{(\mathcal{G},c)\in\pi_{0}\left(\widetilde{\mathsf{LG}}_{g,n,d}(r)\right)}\overline{\mathcal{M}}_{\mathcal{G},c}.$
@@
4.2. Localisation graphs for ¯$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)$
@@
The element $\tau$ is a flip of $C_{k}$ about the perpendicular bisector of the edge between 111 and $k$ and $\rho$ is a rotation. All future references to the dihedral group $D_{k}$ are to this particular subgroup of $B_{k}$.
@@
The hyperoctahedral group $B_{k}$ is essential to the geometric picture. Our computation of $\mathsf{b}_{1,r}^{\mathrm{nrt}}$ sees ¯$\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)$ as gluing caterpillar curves along marked points to a cycle and then decorating the cycle. The wreath product is necessary for the first step as it records the half-edge orientations on each vertex, which is not seen by the standard embedding [^3] Further, the dihedral group fails to be a subgroup of the symmetric group when $k=2$, providing another motivation for the appearance of the wreath product. $D_{k}<S_{k}$. This point of view has been taken by Petersen in [Pet12] for the calculation of $\chi^{S_{n}}(\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}})$.
@@
[…]
@@
Then $\Gamma_{r,d}(k)$ is a set of isomorphism class representatives for the groupoid $\mathsf{LG}_{1,0,d}^{\circ}(r)^{k}$, and given $[\delta]\in\Gamma_{r,d}(k)$, a choice of lift to $\delta\in\tilde{\Gamma}_{r,d}(k)$ defines a subgroup $D_{k}(\delta)$ of $D_{k}$, by taking the stabiliser of $\delta$. We encode the combinatorics of these decorations as follows: for each integer $r\geq 1$, define a graded $\mathbb{B}$-space $\mathsf{Dih}_{r}$ by
@@ 4.3. Caterpillars
Let $n\geq 1$. The moduli space $\text{Cat}_{2,n}$ of $n$-pointed caterpillars is the locally closed subscheme of ¯$\overline{\mathcal{M}}_{0,n+2}$ consisting of pointed rational stable curves $(C,\{q_{i}\}_{i=1}^{n+2})$ that satisfy the following two conditions:
@@
Let $S_{2,n}\subset S_{n+2}$ be the Young subgroup of the partition $[n+2]=\{1,2\}\sqcup\{3,\dots,n+2\}$, so that $S_{2,n}\cong S_{2}\times S_{n}$. We observe that $\mathrm{Cat}_{2,n}$ is an $S_{2}\times S_{n}$-invariant subspace of ¯$\overline{\mathcal{M}}_{0,n+2}$.
@@
¯$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},0)^{\mathbb{C}^{\star}}\cong\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}\times\left(\mathbb{P}^{r}\right)^{\mathbb{C}^{\star}}.$
@@
¯$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}\cong\coprod_{k\geq 2}\coprod_{[\delta]\in\Gamma_{r,d}(k)}\left(\coprod_{\begin{subarray}{c}{n_{1}+\cdots+n_{k}=n}\\ {n_{i}\geq 0}\end{subarray}}\operatorname{Ind}_{S_{n_{1}}\times\cdots\times S_{n_{k}}}^{S_{n}}\mathsf{Cat}(n_{1})\times\cdots\times\mathsf{Cat}(n_{k})\right)/D_{k}(\delta),$
@@
¯$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}\cong\coprod_{k\geq 2}\coprod_{(\mathcal{G},c)\in\pi_{0}\left(\widetilde{\mathsf{LG}}^{\circ}_{1,n,d}(r)^{k}\right)}\overline{\mathcal{M}}^{\circ}_{\mathcal{G},c},$
@@
¯$\overline{\mathcal{M}}^{\circ}_{\mathcal{G},c}\subset\overline{\mathcal{M}}_{\mathcal{G},c}$
@@
¯$\overline{\mathcal{M}}^{\circ}_{\mathcal{G},c}\cong\left(\operatorname{Ind}_{\prod_{v\in V(\mathcal{G})}S_{n(v)}}^{S_{n}}\prod_{v\in V(\mathcal{G})}\mathsf{Cat}(n(v))\right)/\operatorname{Aut}(\mathcal{G},c),$
@@
by forgetting the marking function. If we let $[\delta]\in\Gamma_{r,d}(k)$ denote the image of $(\mathcal{G},c)$ under this functor and $\delta\in\tilde{\Gamma}_{r,d}(k)$ its chosen lift, we get an identification of the underlying graph of $\mathcal{G}$ with $C_{k}$, which also gives the data of a function $f:V(C_{k})\to\mathbb{Z}_{\geq 0}$ with $\sum_{i=1}^{k}f(i)=n$, by recording how many markings each vertex of $\mathcal{G}$ supports. Let $\mathrm{Part}(C_{k},n)$ denote the set of such functions. The group $D_{k}(\delta)$ acts on $\mathrm{Part}(C_{k},n)$, and our choice of lift $\delta$ induces an identification of $\operatorname{Aut}(\mathcal{G},c)$ with the stabiliser $\mathrm{Stab}_{D_{k}(\delta)}(f)$. If we let $n_{i}:=f(i)$ under the identification of $V(C_{k})$ with $\{1,\ldots,k\}$, we can write
¯$\overline{\mathcal{M}}^{\circ}_{\mathcal{G},c}\cong\left(\operatorname{Ind}_{\mathrm{Stab}_{D_{k}(\delta)}(f)}^{D_{k}(\delta)}\operatorname{Ind}_{S_{f(1)}\times\cdots\times S_{f(k)}}^{S_{n}}\prod_{i=1}^{k}\mathsf{Cat}(f(i))\right)/D_{k}(\delta).$
The fibre over $[\delta]$ is identified with
@@
¯$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)$ $\cong\coprod_{[\delta]\in\Gamma_{r,d}(k)}\coprod_{[f]\in\frac{\mathrm{Part}(C_{k},n)}{D_{k}(\delta)}}\left(\operatorname{Ind}_{\mathrm{Stab}_{D_{k}(\delta)}(f)}^{D_{k}(\delta)}\operatorname{Ind}_{S_{f(1)}\times\cdots\times S_{f(k)}}^{S_{n}}\prod_{i=1}^{k}\mathsf{Cat}(f(i))\right)/D_{k}(\delta)$

### v2

@@ 3.4. Separating rational tails
The term $\operatorname{Spec}\mathbb{C}]$ in the formula has bidegree $n=1,d=0$ and is interpreted as providing legs on the core. In the following proof, it will be convenient to interpret the term by formally setting $\overline{\mathcal{M}}_{0}^{\star}(r)(1,0):=\operatorname{Spec}\mathbb{C}$, whereas $\overline{\mathcal{M}}_{0}^{\star}(r)(0,0)=\varnothing$ because of the stability condition.
@@
Unwrapping the definition of plethysms of graded $\mathbb{S}$-spaces, the assignment of $(n,d)$ on the right hand side is equal to
@@
Denote the item in the above formula indexed by $\delta,j$ as $X_{\delta,j}$. We claim that $X_{\delta,j}$ has the same class as the locally closed stratum $\overline{\mathcal{M}}^{(\delta,j)}_{1,n}(\mathbb{P}^{r},d)$ in $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ consisting of stable maps where the core has degree $\delta$, and the core has valency $j$: in other words, the number of rational tails and the number of legs on the core sum to $j$.
To prove the claim, we further stratify $\overline{\mathcal{M}}^{(\delta,j)}_{1,n}(\mathbb{P}^{r},d)$ by the unordered partitions $\lambda\vdash n$ and $\mu\vdash(d-\delta)$ into $j$ non-negative parts, which record the distribution of the number of marked points and degrees onto each leg or rational tail. Let $\overline{\mathcal{M}}^{(\lambda,\mu)}_{1,n}(\mathbb{P}^{r},d)$ be such a refined stratum. From Lemmas 3.6, 3.7, we have that
$\overline{\mathcal{M}}^{(\lambda,\mu)}_{1,n}(\mathbb{P}^{r},d)=\operatorname{Ind}^{S_{n}}_{S_{\lambda}}\left[\overline{\mathcal{M}}^{\mathrm{nrt}}_{1,j}(\mathbb{P}^{r},\delta)\times\prod_{i=1}^{j}\overline{\mathcal{M}}^{\star}_{0}(r)(\lambda_{i},\mu_{i})\right]/\mathrm{Aut}(\lambda,\mu),$
@@
where each $Y_{(\lambda,\mu)}$ is a union of $j$-fold products in $\left(\mathsf{pt}_{S_{1}}+\overline{\mathcal{M}}_{0}^{*}(r)\right)^{\boxtimes j}(n,d-\delta)$ such that the unordered partitions on marked points and degrees recover $\lambda,\mu$. More explicitly, consider the set $L_{(\lambda,\mu)}$ of pairs of ordered partitions that forget to $(\lambda,\mu)$, then $Y_{(\lambda,\mu)}$ is a disjoint union
@@
Applying Lemma 3.1 to the $S_{j}$-action on $\overline{\mathcal{M}}^{\mathrm{nrt}}_{1,j}(\mathbb{P}^{r},\delta)\times Y_{(\lambda,\mu)}$, we hence identify
$\left[\overline{\mathcal{M}}_{1,j}^{\mathrm{nrt}}(\mathbb{P}^{r},\delta)\times Y_{(\lambda,\mu)}\right]/S_{j}\cong\overline{\mathcal{M}}^{(\lambda,\mu)}_{1,n}(\mathbb{P}^{r},d).$
Notice that all the equalities and isomorphisms above are $S_{n}$-equivariant. Summing over all partitions $(\lambda,\mu)$ and all $(\delta,j)$, we get the desired formula in the equivariant Grothendieck ring. ∎
@@ 4. Localisation on the Kontsevich space
Pick a generic $\mathbb{C}^{\star}$-action on $\mathbb{P}^{r}$ so that its fixed points are $(r+1)$ isolated points. This induces an action of $\mathbb{C}^{\star}$ on the $\mathbb{S}$-space $\overline{\mathcal{M}}_{g}(r)$, which preserves the subspace $\overline{\mathcal{M}}_{g}^{\mathrm{nrt}}(r)$. Lemma 3.2 gives
@@
In this section, we introduce a combinatorial graded $\mathbb{B}$-space $\mathsf{Dih}_{r}$ and an $(S_{2}\times\mathbb{S})$-space $\mathsf{Cat}$ of chains of rational curves, and prove the following theorem for the motive $\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)]\in K_{0}(\mathsf{V},\mathbb{S})[[q]]$.
@@
$\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)^{\mathbb{C}^{\star}}]=[({\mathbb{P}^{r}})^{\mathbb{C}^{\star}}][\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(0)]+[\mathsf{Dih}_{r}]\circ_{S_{2}}[\mathsf{Cat}].$
Graber and Pandharipande [31, §4] have described the components
$\overline{\mathcal{M}}_{g}(r)^{\mathbb{C}^{\star}}(n,d)=\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$
of the graded $\mathbb{S}$-space $\overline{\mathcal{M}}_{g}(r)^{\mathbb{C}^{\star}}$ in terms of certain decorated graphs, as we recall shortly. Theorem 4.1 is a motivic interpretation of their description when $g=1$.
@@ 4.1. Localisation graphs
is a graph colouring of $\mathbf{G}$ with $r+1$ colours. The isomorphisms in $\mathsf{LG}_{g,n,d}(r)$ are $\mathsf{LG}_{g,n,d}$-isomorphisms that also respect the colouring. For convenience, we set $\mathsf{LG}_{g,n,d}(0):=\mathsf{LG}_{g,n,d}$. An object $(\mathbf{G},c)$ of the groupoid $\mathsf{LG}_{g,n,d}(r)$ corresponds to a $\mathbb{C}^{\star}$-fixed closed subscheme $\overline{\mathcal{M}}_{\mathbf{G},c}\subset\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ with
(4.2) $\overline{\mathcal{M}}_{\mathbf{G},c}\cong\prod_{v\in V(\mathbf{G})}\overline{\mathcal{M}}_{w(v),n(v)+\mathrm{val}(v)}/\mathrm{Aut}_{\mathsf{LG}_{g,n,d}(r)}(\mathbf{G},c),$
@@
(4.3) $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}=\coprod_{(\mathbf{G},c)\in\pi_{0}(\mathsf{LG}_{g,n,d}(r))}\overline{\mathcal{M}}_{\mathbf{G},c},$
@@
Then $\operatorname{Aut}(\mathcal{G},c)$ is a subgroup of $\operatorname{Aut}(\mathcal{G})$. Each $(\mathcal{G},c)$ in the groupoid $\widetilde{\mathsf{LG}}_{g,n,d}(r)$ defines an $S_{n}$-equivariant $\mathbb{C}^{\star}$-fixed subscheme $\overline{\mathcal{M}}_{\mathcal{G},c}$ of $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$, leading to $S_{n}$-equivariant versions of equations (4.2) and (4.3):
(4.4) $\overline{\mathcal{M}}_{\mathcal{G},c}\cong\left(\operatorname{Ind}_{\prod_{v\in V(\mathcal{G})}S_{n(v)}}^{S_{n}}\prod_{v\in V(\mathcal{G})}\overline{\mathcal{M}}_{g(v),n(v)+\operatorname{val}(v)}\right)/\operatorname{Aut}(\mathcal{G},c),$
(4.5) $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}=\coprod_{(\mathcal{G},c)\in\pi_{0}\left(\widetilde{\mathsf{LG}}_{g,n,d}(r)\right)}\overline{\mathcal{M}}_{\mathcal{G},c}.$
@@
4.2. Localisation graphs for $\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)$
@@
The element $\tau$ is a flip of $C_{k}$ about the perpendicular bisector of the edge between $1$ and $k$ and $\rho$ is a rotation. All future references to the dihedral group $D_{k}$ are to this particular subgroup of $B_{k}$.
@@
The hyperoctahedral group $B_{k}$ is essential to the geometric picture. Our computation of $\mathsf{b}_{1,r}^{\mathrm{nrt}}$ sees $\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)$ as gluing caterpillar curves along marked points to a cycle and then decorating the cycle. The wreath product is necessary for the first step as it records the half-edge orientations on each vertex, which is not seen by the standard embedding [^3] Further, the dihedral group fails to be a subgroup of the symmetric group when $k=2$, providing another motivation for the appearance of the wreath product. $D_{k}<S_{k}$. This point of view has been taken by Petersen in [54] for the calculation of $\chi^{S_{n}}(\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}})$.
@@
The definition of $(r,d)$-decorations makes sense on an arbitrary graph. In the case of $C_{k}$, an $(r,d)$-decoration defines an object of the groupoid $\mathsf{LG}_{1,0,d}(r)^{k}$.
@@
Then $\Gamma_{r,d}(k)$ is a set of isomorphism class representatives for the groupoid $\mathsf{LG}_{1,0,d}^{\circ}(r)^{k}$, and given $\delta]\in\Gamma_{r,d}(k)$, a choice of lift to $\delta\in\tilde{\Gamma}_{r,d}(k)$ defines a subgroup $D_{k}(\delta)$ of $D_{k}$, by taking the stabiliser of $\delta$. We encode the combinatorics of these decorations as follows: for each integer $r\geq 1$, define a graded $\mathbb{B}$-space $\mathsf{Dih}_{r}$ by
@@ 4.3. Caterpillars
Let $n\geq 1$. The moduli space $\text{Cat}_{2,n}$ of $n$-pointed caterpillars is the locally closed subscheme of $\overline{\mathcal{M}}_{0,n+2}$ consisting of pointed rational stable curves $(C,\{q_{i}\}_{i=1}^{n+2})$ that satisfy the following two conditions:
@@
Let $S_{2,n}\subset S_{n+2}$ be the Young subgroup of the partition $n+2]=\{1,2\}\sqcup\{3,\dots,n+2\}$, so that $S_{2,n}\cong S_{2}\times S_{n}$. We observe that $\mathrm{Cat}_{2,n}$ is an $S_{2}\times S_{n}$-invariant subspace of $\overline{\mathcal{M}}_{0,n+2}$.
@@
$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},0)^{\mathbb{C}^{\star}}\cong\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}\times\left(\mathbb{P}^{r}\right)^{\mathbb{C}^{\star}}.$
@@
$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}\cong\coprod_{k\geq 2}\coprod_{[\delta]\in\Gamma_{r,d}(k)}\left(\coprod_{\begin{subarray}{c}{n_{1}+\cdots+n_{k}=n}\\ {n_{i}\geq 0}\end{subarray}}\operatorname{Ind}_{S_{n_{1}}\times\cdots\times S_{n_{k}}}^{S_{n}}\mathsf{Cat}(n_{1})\times\cdots\times\mathsf{Cat}(n_{k})\right)/D_{k}(\delta),$
@@
$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}\cong\coprod_{k\geq 2}\coprod_{(\mathcal{G},c)\in\pi_{0}\left(\widetilde{\mathsf{LG}}^{\circ}_{1,n,d}(r)^{k}\right)}\overline{\mathcal{M}}^{\circ}_{\mathcal{G},c},$
@@
$\overline{\mathcal{M}}^{\circ}_{\mathcal{G},c}\subset\overline{\mathcal{M}}_{\mathcal{G},c}$
@@
$\overline{\mathcal{M}}^{\circ}_{\mathcal{G},c}\cong\left(\operatorname{Ind}_{\prod_{v\in V(\mathcal{G})}S_{n(v)}}^{S_{n}}\prod_{v\in V(\mathcal{G})}\mathsf{Cat}(n(v))\right)/\operatorname{Aut}(\mathcal{G},c),$
@@
by forgetting the marking function. If we let $\delta]\in\Gamma_{r,d}(k)$ denote the image of $(\mathcal{G},c)$ under this functor and $\delta\in\tilde{\Gamma}_{r,d}(k)$ its chosen lift, we get an identification of the underlying graph of $\mathcal{G}$ with $C_{k}$, which also gives the data of a function $f:V(C_{k})\to\mathbb{Z}_{\geq 0}$ with $\sum_{i=1}^{k}f(i)=n$, by recording how many markings each vertex of $\mathcal{G}$ supports. Let $\mathrm{Part}(C_{k},n)$ denote the set of such functions. The group $D_{k}(\delta)$ acts on $\mathrm{Part}(C_{k},n)$, and our choice of lift $\delta$ induces an identification of $\operatorname{Aut}(\mathcal{G},c)$ with the stabiliser $\mathrm{Stab}_{D_{k}(\delta)}(f)$. If we let $n_{i}:=f(i)$ under the identification of $V(C_{k})$ with $\{1,\ldots,k\}$, we can write
$\overline{\mathcal{M}}^{\circ}_{\mathcal{G},c}\cong\left(\operatorname{Ind}_{\mathrm{Stab}_{D_{k}(\delta)}(f)}^{D_{k}(\delta)}\operatorname{Ind}_{S_{f(1)}\times\cdots\times S_{f(k)}}^{S_{n}}\prod_{i=1}^{k}\mathsf{Cat}(f(i))\right)/D_{k}(\delta).$
The fibre over $\delta]$ is identified with
@@
$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)$ $\cong\coprod_{[\delta]\in\Gamma_{r,d}(k)}\coprod_{[f]\in\frac{\mathrm{Part}(C_{k},n)}{D_{k}(\delta)}}\left(\operatorname{Ind}_{\mathrm{Stab}_{D_{k}(\delta)}(f)}^{D_{k}(\delta)}\operatorname{Ind}_{S_{f(1)}\times\cdots\times S_{f(k)}}^{S_{n}}\prod_{i=1}^{k}\mathsf{Cat}(f(i))\right)/D_{k}(\delta)$